If the perimeter of a rectangular courtyard is 36 yd, and the length is 3 yd longer than twice the width, then what are the length and width?
Answer by mananth(16949)

If the perimeter of a rectangular courtyard is 36 yd, and the length is 3 yd longer than twice the width, then what are the length and width?
..
widthe be x
length = 2x+3
..
Perimeter = 2L+2W
2x+2(2x+3)=36
2x+4x+6=36
6x+6=36
6x=30
x=5
width = 5
length = 2*5 +3=13
..
CHECK
2*13+2*5=36